# A Survey of Vibe Coding with Large Language Models

### v1

@@ 7.4.2 Multi-Agent Collaborative Feedback
Multi-agent systems employ several distinct feedback mechanisms enabling collaborative code improvement [194, 111] with most fundamental distinctions lying between inter-agent feedback and self-feedback mechanisms, with inter-agent feedback involving agents providing constructive criticism to each other based on interactions and collaborations helping agents identify improvement areas and adapt strategies for continuous learning within systems [195, 10, 127]. In multi-core agent systems, intra-agent feedback emerges as agents exchange information and share observations including shared observations, alternative perspectives on problems, and evaluations of proposed plans promoting collaborative problem-solving and cross-validation enhancing overall system robustness through collective intelligence [196]. Self-feedback mechanisms allow agents to assess own performance and identify improvement areas through self-assessment based on predefined criteria or goals, while external feedback-based optimization leverages evaluative signals from external models or agents to refine agent behavior integrating external reflections and corrections to enhance robustness and adaptability [195].
Core collaborative feedback mechanisms center on peer-reflection involving information exchange through role specialization and structured communication. Self-feedback loops represent another key collaborative mechanism where agents iteratively refine strategies based on both quantitative numerical evaluations and qualitative natural language assessments [197]. Feedback-refine cycles enable agents to store language-level memories and collaboratively iterate on solutions with parallel agents exchanging critiques [115]. Mutual assessment and pruning mechanisms facilitate collaborative optimization where agents produce multiple candidates and evaluate outputs from other agents [243].
Self-refinement mechanisms in multi-agent coding systems operate through iterative feedback-refine cycles demonstrating that when run in parallel with agents exchanging critiques feedback-refine cycles significantly boost code generation quality [115]. Hybrid value functions represent key advancements incorporating both quantitative and qualitative assessment mechanisms [197]. Circular optimization mechanisms create continuous improvement loops through self-reflection scoring systems with notable frameworks like CodeCoR introducing reflection agents between code generation, testing, and repair stages [243, 127].
@@ 7.4.3 Reflection and Memory-Based Feedback
Several foundational frameworks have emerged as core approaches for implementing self-refinement in large language models with most notable being Self-Refine which introduced comprehensive end-to-end self-correction approaches [198, 105, 547, 199]. Reflexion represents another foundational framework using verbal reinforcement to help agents learn from prior failures incorporating three key LLM-based modules: Actors that generate actions, Evaluators that assess results, and Self-Reflection models that analyze failure cases [201, 199, 548].
Memory architecture represents critical components in self-reflection systems with most frameworks adopting dual-memory approaches separating short-term and long-term memory components, with trajectory history serving as short-term memory while outputs from Self-Reflection models are stored in long-term memory [201, 79, 272, 200, 90, 549]. Transformation of environmental feedback into memory representations is key innovation with Self-Reflection models taking sparse reward signals along with current trajectory and persistent memory then generating nuanced and specific feedback, effectively converting scalar rewards into natural language feedback [201, 547, 105].
Coding domain has emerged as particularly fertile ground for self-refinement applications with numerous frameworks specifically designed to leverage unique feedback mechanisms with program execution feedback representing most common approaches where frameworks like Self-Edit and Self-Evolve execute initial programs on test cases and provide execution results back as feedback [550]. Multi-agent debugging architectures have also gained prominence with frameworks like Refinement and Guidance Debugging (RGD) system introducing multi-LLM-based agent debuggers [118].
Performance evaluation of self-refinement frameworks reveals significant advances with Reflexion achieving particularly notable results demonstrating 91% pass@1 accuracy on HumanEval coding benchmark surpassing previous GPT-4 performance of 80% representing 11 percentage point improvements achieved through verbal reinforcement [201]. Self-Refine framework has demonstrated consistent performance gains achieving approximately 20% absolute improvement in task performance across diverse applications while requiring minimal computational resources [548].
@@ 8 Vibe Coding Development Models
Figure 8: Comparison of Vibe Coding Development Models Framework With Software Development Models [551, 552, 553].
@@ 8.2 Unconstrained Automation Model (UAM)
This model represents the development approach most closely aligned with the original definition of Vibe Coding, characterized by complete trust in AI output, minimal code scrutiny, and correctness validation through functionality testing rather than code comprehension. Whether developers issue short instructions through high-frequency dialogues or allow AI agents to autonomously decompose tasks and iterate multiple rounds, any workflow that accepts output without deep code inspection falls within this paradigm. This model emphasizes high development velocity and low technical barriers, enabling non-programmers to rapidly construct application prototypes, with development timelines compressed to several times or even dozens of times faster than traditional approaches. This paradigm bears conceptual similarity to Rapid Application Development (RAD) methodologies in traditional software engineering, which similarly emphasize rapid prototyping and shortened development cycles over extensive upfront planning [554, 553]. Both approaches prioritize speed and iterative refinement, though UAM relies on AI automation whereas RAD depends on specialized development tools and frameworks [273].
@@ 8.3 Iterative Conversational Collaboration Model (ICCM)
ICCM positions AI as a programming partner rather than a fully autonomous agent, constructing software through continuous dialogue and iterative cycles while humans maintain comprehensive oversight, understanding, and control over code quality. This model requires developers to review and comprehend each AI output before deciding whether to accept it, forming an iterative workflow of “AI generates $\rightarrow$ human reviews and understands $\rightarrow$ testing validates $\rightarrow$ acceptance decision.” This approach parallels pair programming in agile development [555, 156], where AI assumes the Driver role by rapidly producing code while humans serve as the Navigator, guiding direction and reviewing quality [556]. Empirical studies on pair programming have demonstrated improvements in code quality and knowledge transfer [557, 558], benefits that similarly manifest in ICCM through the continuous human-AI collaboration loop. The iterative nature manifests in continuous improvement and incremental refinement across multiple dialogue rounds.
@@ 8.4 Planning-Driven Model (PDM)
PDM emphasizes establishing clear development plans and design specifications by humans before AI coding commences, then guiding AI to implement progressively according to the plan, applying the “architecture-first” philosophy of traditional software engineering to Vibe Coding. Developers invest time in organizing module decomposition, data structures, and feature lists, typically producing a three-document system comprising technical specification documents, coding rule files, and implementation plans, which serve as the “blueprint” for AI coding. This model conceptually aligns with the waterfall model in traditional software development [551, 273], where sequential phases—requirements analysis, design, implementation, and testing—proceed in a structured, linear fashion. However, PDM retains greater flexibility than traditional waterfall approaches, as AI can rapidly iterate on implementations while adhering to the established architectural framework [482]. This model parallels the traditional development division where architects design while teams implement, except here the “team” is undertaken by AI, with humans playing the roles of architect and project manager.
@@ 8.5 Test-Driven Model (TDM)
TDM applies the TDD philosophy from traditional software engineering to Vibe Coding, with the core idea of first defining tests and acceptance criteria, having AI generate code satisfying the tests, and ensuring code quality through automated testing rather than manual review. Unlike other models that constrain AI behavior through natural language descriptions or documentation, TDM employs test cases as precise constraints on AI, where tests serve as specifications, explicitly defining objective success criteria. This model directly inherits from TDD practices established in software engineering [529, 156], following the classic red-green-refactor cycle: write failing tests, implement code to pass tests, then refactor for quality [274]. Humans write test cases, AI writes implementation code, iterating development through the classic TDD cycle.
@@ 9.1 Reengineering of Development Process in Vibe Coding
The central premise of Vibe Coding—validating implementation through outcome observation rather than line-by-line comprehension—introduces notable shifts in software development practices. Traditional Software Development Life Cycle (SDLC) models, from Waterfall to Scrum [139], were designed around assumptions of human-authored code and predictable progression through discrete phases. Vibe Coding extends these models by introducing AI-mediated development cycles that operate at accelerated timescales, prompting adaptations in process design, expanding developer responsibilities, and surfacing new considerations in project management and quality assurance. While empirical research on these impacts remains nascent and lacks longitudinal validation [559], we can synthesize characteristics of this emergent paradigm [560, 20] to project its evolutionary trajectory within existing development frameworks.
@@ 9.1.1 From Phased Lifecycles to Continuous Micro-Iterations
Building upon the traditional "edit-compile-debug" loop, Vibe Coding introduces a complementary "prompt-generate-validate" cycle that operates at compressed timescales—from the weeks typical of Agile sprints to seconds or minutes—thereby enabling more rapid iteration within the existing development framework. This accelerated cycle is organized around "iterative goal satisfaction cycles" [560], where the developer’s workflow becomes a continuous conversation with the coding agent, and each exchange generates a testable artifact [561, 562]. This blurs the traditional boundaries between SDLC phases [562].
@@
Instead of creating comprehensive design documents upfront, a developer can engage in exploratory Vibe Coding [563]. The initial "vibe" might be a high-level user story, which the developer translates into a series of prompts to generate a functional prototype. The act of interacting with this AI-generated prototype refines the requirements in real-time [561]. The design emerges organically from this iterative dialogue, transforming the design phase from a distinct, pre-implementation step into a continuous, integrated activity.
@@
In this paradigm, the traditional implementation phase undergoes significant transformation, with coding agents handling much of the syntactic construction while developers focus on higher-level orchestration. The developer’s role evolves to encompass both traditional code authorship and emergent responsibilities as system director and prompt engineer, with the balance tilting toward orchestration as agent capabilities advance [564, 243]. While syntactic mastery remains foundational, the skill portfolio increasingly emphasizes the ability to articulate intent, manage context, and evaluate agent output [565]—representing an expansion rather than replacement of traditional programming competencies [560].
@@
The developer’s primary "vibe check" is a form of continuous, informal acceptance testing. However, this shift introduces ambiguity. While this immediate feedback is powerful for rapid iteration, it lacks the rigor of formal Quality Assurance [566, 567]. The challenge lies in integrating structured testing methodologies, such as the Test-Driven model [529] mentioned in our taxonomy, where the developer first defines formal tests that the AI agent must then generate code to pass. This provides a concrete, verifiable "vibe" for the agent to target.
@@ 9.1.2 Redefinition of Developer Roles and Skillsets
Among these emergent capabilities, the ability to translate complex requirements into effective prompts has proven particularly valuable in practice [20]. This is a creative and analytical skill, requiring an understanding of how the model "thinks" without needing to understand its internal architecture.
@@
When AI-generated systems fail, traditional line-by-line debugging becomes less effective, as the generated logic can be opaque or non-intuitive [568]. Debugging evolves toward a process of hypothesizing system-level failures, isolating problematic components through targeted tests, and then using prompts to guide the AI to regenerate or patch the faulty section [569]. The focus shifts from algorithmic debugging to behavioral debugging.
@@
As identified in our analysis, successful Vibe Coding depends heavily on context engineering. Successful practitioners develop expertise in context curation, learning to feed agents the right information—APIs, data schemas, existing codebase snippets, design patterns—to constrain generation space effectively. The current ecosystem lacks mature open-source or commercial platforms specifically designed for this dynamic context management in coding workflows [560, 570], making it a largely manual and ad-hoc process.
@@ 9.1.3 New Challenges in Project Management and Collaboration
The Vibe Coding paradigm introduces significant hurdles for traditional project management. Estimating effort for a task becomes difficult; a complex feature might be generated in minutes with the right prompt, while a seemingly simple one could take hours of iterative refinement [571]. Code reviews, a cornerstone of traditional quality control [572, 573], face new challenges in this context. When reviewing AI-generated code that the original prompter does not fully understand, the review focus naturally expands beyond line-level correctness to also include validation of prompt history, generated tests, and observed behavior against requirements [574, 575, 576].
Team collaboration also changes, with "pair programming" [577] potentially evolving into "mob prompting," [578] where multiple developers collaborate on crafting the prompts and context for a shared AI agent. However, the lack of documented case studies of Vibe Coding in large-scale enterprise projects means these process changes remain largely theoretical and "under-explored" [579, 580, 560, 581, 582, 583] representing a critical area for future empirical research.

### v2

@@ 7.4.2 Multi-Agent Collaborative Feedback
Multi-agent systems employ several distinct feedback mechanisms enabling collaborative code improvement [ishibashi2024self, ashrafi2025enhancing] with most fundamental distinctions lying between inter-agent feedback and self-feedback mechanisms, with inter-agent feedback involving agents providing constructive criticism to each other based on interactions and collaborations helping agents identify improvement areas and adapt strategies for continuous learning within systems [talebirad2023multi, qian2023chatdev, islam2024mapcoder]. In multi-core agent systems, intra-agent feedback emerges as agents exchange information and share observations including shared observations, alternative perspectives on problems, and evaluations of proposed plans promoting collaborative problem-solving and cross-validation enhancing overall system robustness through collective intelligence [hassouna2024llm]. Self-feedback mechanisms allow agents to assess own performance and identify improvement areas through self-assessment based on predefined criteria or goals, while external feedback-based optimization leverages evaluative signals from external models or agents to refine agent behavior integrating external reflections and corrections to enhance robustness and adaptability [talebirad2023multi].
Core collaborative feedback mechanisms center on peer-reflection involving information exchange through role specialization and structured communication. Self-feedback loops represent another key collaborative mechanism where agents iteratively refine strategies based on both quantitative numerical evaluations and qualitative natural language assessments [antoniades2024swe]. Feedback-refine cycles enable agents to store language-level memories and collaboratively iterate on solutions with parallel agents exchanging critiques [ando2025when]. Mutual assessment and pruning mechanisms facilitate collaborative optimization where agents produce multiple candidates and evaluate outputs from other agents [dong2025survey].
Self-refinement mechanisms in multi-agent coding systems operate through iterative feedback-refine cycles demonstrating that when run in parallel with agents exchanging critiques feedback-refine cycles significantly boost code generation quality [ando2025when]. Hybrid value functions represent key advancements incorporating both quantitative and qualitative assessment mechanisms [antoniades2024swe]. Circular optimization mechanisms create continuous improvement loops through self-reflection scoring systems with notable frameworks like CodeCoR introducing reflection agents between code generation, testing, and repair stages [dong2025survey, islam2024mapcoder].
@@ 7.4.3 Reflection and Memory-Based Feedback
Several foundational frameworks have emerged as core approaches for implementing self-refinement in large language models with most notable being Self-Refine which introduced comprehensive end-to-end self-correction approaches [mousavi2023critics, madaan2023self, chang2024efficient, park2025self]. Reflexion represents another foundational framework using verbal reinforcement to help agents learn from prior failures incorporating three key LLM-based modules: Actors that generate actions, Evaluators that assess results, and Self-Reflection models that analyze failure cases [shinn2023reflexion, park2025self, kamalov2025evolution].
Memory architecture represents critical components in self-reflection systems with most frameworks adopting dual-memory approaches separating short-term and long-term memory components, with trajectory history serving as short-term memory while outputs from Self-Reflection models are stored in long-term memory [shinn2023reflexion, yao2023tree, dong2023self, ge2025samule, park2023generative, liu2024roleagent]. Transformation of environmental feedback into memory representations is key innovation with Self-Reflection models taking sparse reward signals along with current trajectory and persistent memory then generating nuanced and specific feedback, effectively converting scalar rewards into natural language feedback [shinn2023reflexion, chang2024efficient, madaan2023self].
Coding domain has emerged as particularly fertile ground for self-refinement applications with numerous frameworks specifically designed to leverage unique feedback mechanisms with program execution feedback representing most common approaches where frameworks like Self-Edit and Self-Evolve execute initial programs on test cases and provide execution results back as feedback [pan2023automatically]. Multi-agent debugging architectures have also gained prominence with frameworks like Refinement and Guidance Debugging (RGD) system introducing multi-LLM-based agent debuggers [jin2024rgd].
Performance evaluation of self-refinement frameworks reveals significant advances with Reflexion achieving particularly notable results demonstrating 91% pass@1 accuracy on HumanEval coding benchmark surpassing previous GPT-4 performance of 80% representing 11 percentage point improvements achieved through verbal reinforcement [shinn2023reflexion]. Self-Refine framework has demonstrated consistent performance gains achieving approximately 20% absolute improvement in task performance across diverse applications while requiring minimal computational resources [kamalov2025evolution].
@@ 8 Vibe Coding Development Models
Figure 8: Comparison of Vibe Coding Development Models Framework With Software Development Models [royce1987managing, manifesto2001manifesto, martin1991rapid].
@@ 8.2 Unconstrained Automation Model (UAM)
This model represents the development approach most closely aligned with the original definition of Vibe Coding, characterized by complete trust in AI output, minimal code scrutiny, and correctness validation through functionality testing rather than code comprehension. Whether developers issue short instructions through high-frequency dialogues or allow AI agents to autonomously decompose tasks and iterate multiple rounds, any workflow that accepts output without deep code inspection falls within this paradigm. This model emphasizes high development velocity and low technical barriers, enabling non-programmers to rapidly construct application prototypes, with development timelines compressed to several times or even dozens of times faster than traditional approaches. This paradigm bears conceptual similarity to Rapid Application Development (RAD) methodologies in traditional software engineering, which similarly emphasize rapid prototyping and shortened development cycles over extensive upfront planning [beynon1999rapid, martin1991rapid]. Both approaches prioritize speed and iterative refinement, though UAM relies on AI automation whereas RAD depends on specialized development tools and frameworks [dima2018waterfall].
@@ 8.3 Iterative Conversational Collaboration Model (ICCM)
ICCM positions AI as a programming partner rather than a fully autonomous agent, constructing software through continuous dialogue and iterative cycles while humans maintain comprehensive oversight, understanding, and control over code quality. This model requires developers to review and comprehend each AI output before deciding whether to accept it, forming an iterative workflow of “AI generates $\rightarrow$ human reviews and understands $\rightarrow$ testing validates $\rightarrow$ acceptance decision.” This approach parallels pair programming in agile development [williams2000strengthening, beck2000extreme], where AI assumes the Driver role by rapidly producing code while humans serve as the Navigator, guiding direction and reviewing quality [begel2008pair]. Empirical studies on pair programming have demonstrated improvements in code quality and knowledge transfer [hannay2009effectiveness, bipp2008pair], benefits that similarly manifest in ICCM through the continuous human-AI collaboration loop. The iterative nature manifests in continuous improvement and incremental refinement across multiple dialogue rounds.
@@ 8.4 Planning-Driven Model (PDM)
PDM emphasizes establishing clear development plans and design specifications by humans before AI coding commences, then guiding AI to implement progressively according to the plan, applying the “architecture-first” philosophy of traditional software engineering to Vibe Coding. Developers invest time in organizing module decomposition, data structures, and feature lists, typically producing a three-document system comprising technical specification documents, coding rule files, and implementation plans, which serve as the “blueprint” for AI coding. This model conceptually aligns with the waterfall model in traditional software development [royce1987managing, dima2018waterfall], where sequential phases—requirements analysis, design, implementation, and testing—proceed in a structured, linear fashion. However, PDM retains greater flexibility than traditional waterfall approaches, as AI can rapidly iterate on implementations while adhering to the established architectural framework [highsmith2002agile]. This model parallels the traditional development division where architects design while teams implement, except here the “team” is undertaken by AI, with humans playing the roles of architect and project manager.
@@ 8.5 Test-Driven Model (TDM)
TDM applies the TDD philosophy from traditional software engineering to Vibe Coding, with the core idea of first defining tests and acceptance criteria, having AI generate code satisfying the tests, and ensuring code quality through automated testing rather than manual review. Unlike other models that constrain AI behavior through natural language descriptions or documentation, TDM employs test cases as precise constraints on AI, where tests serve as specifications, explicitly defining objective success criteria. This model directly inherits from TDD practices established in software engineering [beck2003test, beck2000extreme], following the classic red-green-refactor cycle: write failing tests, implement code to pass tests, then refactor for quality [Janzen2005TestdrivenDC]. Humans write test cases, AI writes implementation code, iterating development through the classic TDD cycle.
@@ 9.1 Reengineering of Development Process in Vibe Coding
The central premise of Vibe Coding—validating implementation through outcome observation rather than line-by-line comprehension—introduces notable shifts in software development practices. Traditional Software Development Life Cycle (SDLC) models, from Waterfall to Scrum [gajbhiye2024secure], were designed around assumptions of human-authored code and predictable progression through discrete phases. Vibe Coding extends these models by introducing AI-mediated development cycles that operate at accelerated timescales, prompting adaptations in process design, expanding developer responsibilities, and surfacing new considerations in project management and quality assurance. While empirical research on these impacts remains nascent and lacks longitudinal validation [Ford2020PandemicDev], we can synthesize characteristics of this emergent paradigm [Sarkar2025VibeCoding, sapkota2025vibe] to project its evolutionary trajectory within existing development frameworks.
@@ 9.1.1 From Phased Lifecycles to Continuous Micro-Iterations
Building upon the traditional "edit-compile-debug" loop, Vibe Coding introduces a complementary "prompt-generate-validate" cycle that operates at compressed timescales—from the weeks typical of Agile sprints to seconds or minutes—thereby enabling more rapid iteration within the existing development framework. This accelerated cycle is organized around "iterative goal satisfaction cycles" [Sarkar2025VibeCoding], where the developer’s workflow becomes a continuous conversation with the coding agent, and each exchange generates a testable artifact [hymel2024ai, wang2025software]. This blurs the traditional boundaries between SDLC phases [wang2025software].
@@
Instead of creating comprehensive design documents upfront, a developer can engage in exploratory Vibe Coding [ulfsnes2024transforming]. The initial "vibe" might be a high-level user story, which the developer translates into a series of prompts to generate a functional prototype. The act of interacting with this AI-generated prototype refines the requirements in real-time [hymel2024ai]. The design emerges organically from this iterative dialogue, transforming the design phase from a distinct, pre-implementation step into a continuous, integrated activity.
@@
In this paradigm, the traditional implementation phase undergoes significant transformation, with coding agents handling much of the syntactic construction while developers focus on higher-level orchestration. The developer’s role evolves to encompass both traditional code authorship and emergent responsibilities as system director and prompt engineer, with the balance tilting toward orchestration as agent capabilities advance [Chen2025ScreenReader, dong2025survey]. While syntactic mastery remains foundational, the skill portfolio increasingly emphasizes the ability to articulate intent, manage context, and evaluate agent output [russo2024navigating]—representing an expansion rather than replacement of traditional programming competencies [Sarkar2025VibeCoding].
@@
The developer’s primary "vibe check" is a form of continuous, informal acceptance testing. However, this shift introduces ambiguity. While this immediate feedback is powerful for rapid iteration, it lacks the rigor of formal Quality Assurance [ferdowsi2024validating, garousi2024ai]. The challenge lies in integrating structured testing methodologies, such as the Test-Driven model [beck2003test] mentioned in our taxonomy, where the developer first defines formal tests that the AI agent must then generate code to pass. This provides a concrete, verifiable "vibe" for the agent to target.
@@ 9.1.2 Redefinition of Developer Roles and Skillsets
Among these emergent capabilities, the ability to translate complex requirements into effective prompts has proven particularly valuable in practice [sapkota2025vibe]. This is a creative and analytical skill, requiring an understanding of how the model "thinks" without needing to understand its internal architecture.
@@
When AI-generated systems fail, traditional line-by-line debugging becomes less effective, as the generated logic can be opaque or non-intuitive [torka2024optimizing]. Debugging evolves toward a process of hypothesizing system-level failures, isolating problematic components through targeted tests, and then using prompts to guide the AI to regenerate or patch the faulty section [vaithilingam2022expectation]. The focus shifts from algorithmic debugging to behavioral debugging.
@@
As identified in our analysis, successful Vibe Coding depends heavily on context engineering. Successful practitioners develop expertise in context curation, learning to feed agents the right information—APIs, data schemas, existing codebase snippets, design patterns—to constrain generation space effectively. The current ecosystem lacks mature open-source or commercial platforms specifically designed for this dynamic context management in coding workflows [Sarkar2025VibeCoding, Dong2018OpenArchitecture], making it a largely manual and ad-hoc process.
@@ 9.1.3 New Challenges in Project Management and Collaboration
The Vibe Coding paradigm introduces significant hurdles for traditional project management. Estimating effort for a task becomes difficult; a complex feature might be generated in minutes with the right prompt, while a seemingly simple one could take hours of iterative refinement [pospieszny2018effective]. Code reviews, a cornerstone of traditional quality control [rigby2013convergent, Ackerman1989SoftwareInspections], face new challenges in this context. When reviewing AI-generated code that the original prompter does not fully understand, the review focus naturally expands beyond line-level correctness to also include validation of prompt history, generated tests, and observed behavior against requirements [sun2025does, rasheed2024ai, arabzadeh2025benchmarking].
Team collaboration also changes, with "pair programming" [zieris2022understanding] potentially evolving into "mob prompting," [buchan2018leveraging] where multiple developers collaborate on crafting the prompts and context for a shared AI agent. However, the lack of documented case studies of Vibe Coding in large-scale enterprise projects means these process changes remain largely theoretical and "under-explored" [Post2005FalconCase, Gadde2025Democratizing, Sarkar2025VibeCoding, watanabe2025use, pandey2024transforming, amasanti2025impact] representing a critical area for future empirical research.